help me solve for y
(x-1)^2 + y^2 = 1

Answer by AnlytcPhil(1810)

help me solve for y
`
` ` ` ` (x-1)2 + y2 = 1
`
Subtract (x-1)2 from both sides
`
` ` ` ` `y2 = 1 - (x-1)2
Write 1 as 12
`
` ` ` ` `y2 = 12 - (x-1)2
`
The right side will factor as the difference of two squares:
`
` ` ` ` `y2 = [1 - (x-1)][1 + (x-1)]
`
Remove inner parentheses
`
` ` ` ` `y2 = [1 - x + 1)(1 + x - 1]
`
Combine like terms
`
` ` ` ` `y2 = [2 - x][x]
`
Rewrite with short factor first
`
` ` ` ` `y2 = x(2-x)
`
Take square roots of both sides, remembering ± on the right
` ` ` ` ` ` ` ` ______
` ` ` ` ` y = ±Öx(2-x)
`
Edwin